a disc of mas 5kg and radius 40 cm rotates about an axis passing through centre and perpendicular to plane.it makes 10 revelutions per second .find angular momentum. in this qo,we don’t get the ans with the formula L=mvr.why [we can find v from v=r omega]

Grade:11

1 Answers

Omesh Nichat
5 years ago
Moment of inertia of a disc=mr^2/2
So by putting values we get, I=2/5kg m^2
No. Of revolutions per second =10×2π=20π rad per second
L=I×omega
   =2/5×20π
    =8π kgm^2/second